OpenStudy (anonymous):

2. Write the equation of the line which passes through (2, 1) and is perpendicular to x = –2.

OpenStudy (anonymous):

Use the equation \[\huge{y=mx+c}\]

OpenStudy (anonymous):

\[\large{x=2}\]\[\large{y=1}\]

OpenStudy (anonymous):

There's no slope here...

OpenStudy (anonymous):

So what would I put for the answer?

OpenStudy (anonymous):

So the equation for the first is y=x-1..So you could apply x and find the other.. :)

OpenStudy (anonymous):

soo... y=-2-1?

OpenStudy (anonymous):

Sorry I went for dinner.. :D As the question asks only for the equation it's y=x-1..

OpenStudy (anonymous):

@mallorymae